# Efficient Quantization-Aware Neural Receivers: Beyond Post-Training Quantization

### v2

@@ 1 Introduction
Deep Learning has recently delivered strong gains at the PHY, where fully-convolutional neural receivers such as DeepRx replace channel estimation, equalization, and demapping to improve error-rate performance under standardized channels and mobility [1]. End-to-end designs further demonstrate that learned receivers can reduce or even eliminate pilot overhead while maintaining reliability, highlighting the practical promise of Artificial Intelligence (AI)-native PHY stacks for beyond-5G and 6G systems [2]. However, deployment in edge-constrained radio hardware remains limited by compute, memory, and latency budgets, making precision reduction via quantization central to practical neural receiver inference. In this context, QAT adapts the model during training to low-precision arithmetic [3, 4, 5, 6, 7, 8, 9].
PTQ avoids retraining and compresses models by up to $8\times$, but can degrade at aggressive bit-widths and under distribution shifts characteristic of high‑Doppler channels. In prior work, a CNN‑based neural receiver with symmetric uniform PTQ (per‑tensor and per‑channel) achieved near‑float32 BLER at 8‑bit and competitive results at 4‑bit, establishing a strong baseline for efficient PHY inference and motivating robustness analysis at low precision [10, 11, 12, 13]. In this paper, we examine whether QAT closes the remaining gap by training with simulated quantizers, thereby improving the dependability of ultra‑low‑bit neural receivers across mobility and channel conditions.
@@ 2.1 Training Objective Under Simulated Quantization
The training objective in QAT is to minimize the task loss when model weights operate under quantized precision while maintaining full-precision activations. The neural receiver function $f(\mathbf{s};W_{q})$ denotes the forward computation that processes an input signal $\mathbf{s}$ using quantized weights $W_{q}=F_{b}(W;\phi_{w})$ to produce the network prediction. Using the fake quantization operator $F_{b}(\cdot)$ defined in (3), the optimization problem becomes:
@@
where $W$ denotes the network weights, $\phi_{w}$ represents the quantization parameters for weights (clipping bounds $\alpha$ and $\beta$), and $\mathcal{D}$ is the training data distribution with signal samples $\mathbf{s}$ and labels $\mathbf{y}$. Note that in this work, activations remain in full precision throughout the network, allowing the neural receiver to preserve signal fidelity while achieving computational efficiency through weight quantization.
@@ 4 Experimental Results and Discussion
LoS comparisons. The lower panel of Figure 1 illustrates the BLER performance under NLoS conditions using the CDL-D channel model. The 8-bit QAT configuration achieves 10% BLER within an SNR range of 3.94–3.97 dB across mobility scenarios, respectively, outperforming the LS baseline by approximately 2.4–4.8 dB as mobility increases. At velocities ranging from 10 to 35 m/s, QAT 4-bit requires 4.20–4.23 dB to attain 10% BLER, maintaining a performance advantage over LS of approximately 2.19–4.50 dB, depending on the mobility level. Results show that quantization-aware training is essential for reliable performance at ultra-low bitwidths, as the 4-bit QAT neural receiver consistently outperforms the PTQ variant across all UE speeds.
@@
5 Conclusion
We developed and evaluated quantization-aware neural receivers, performing a detailed comparison of QAT and PTQ across diverse channel profiles. We used learned weight clipping via the STE and compared QAT- and PTQ-based neural receivers. The 4-bit QAT neural receiver achieves promising performance at 10% BLER in both LoS and NLoS scenarios, in sharp contrast to the poor performance of 4-bit PTQ. It maintains at most a 0.8 dB SNR gap compared to full-precision receivers, demonstrating robustness to aggressive quantization. By enabling up to an $8\times$ model size reduction over FP32, 4-bit quantization supports lightweight inference on resource-constrained devices, reducing memory requirements and inference latency while directly lowering telecommunication capital expenditures through reduced hardware and deployment costs. The strong LoS performance of 4-bit QAT receivers also positions them as viable candidates for integrated sensing and communication as well as site-specific deployments in 6G.

### v3

@@ 1 Introduction
Deep Learning has recently delivered strong gains at the PHY, where fully-convolutional neural receivers such as DeepRx replace channel estimation, equalization, and demapping to improve error-rate performance under standardized channels and mobility [9]. End-to-end designs further demonstrate that learned receivers can reduce or even eliminate pilot overhead while maintaining reliability, highlighting the practical promise of Artificial Intelligence (AI)-native PHY stacks for beyond-5G and 6G systems [2]. However, deployment in edge-constrained radio hardware remains limited by compute, memory, and latency budgets, making precision reduction via quantization central to practical neural receiver inference. In this context, QAT adapts the model during training to low-precision arithmetic [16, 20, 13, 11, 5, 6].
PTQ avoids retraining and compresses models by up to $8\times$, but can degrade at aggressive bit-widths and under distribution shifts characteristic of high‑Doppler channels. In prior work, a CNN‑based neural receiver with symmetric uniform PTQ (per‑tensor and per‑channel) achieved near‑float32 BLER at 8‑bit and competitive results at 4‑bit, establishing a strong baseline for efficient PHY inference and motivating robustness analysis at low precision [22, 15, 14, 19, 21]. In this paper, we examine whether QAT closes the remaining gap by training with simulated quantizers, thereby improving the dependability of ultra‑low‑bit neural receivers across mobility and channel conditions.
@@ 2.1 Training Objective Under Simulated Quantization
The training objective in QAT is to minimize the task loss when model weights operate under quantized precision while maintaining full-precision activations. The neural receiver function $f(\mathbf{s};W_{q})$ denotes the forward computation that processes an input signal $\mathbf{s}$ using quantized weights $W_{q}=F_{b}(W;\phi_{w})$ to produce the network prediction, where $W$ denotes the network weights, $\phi_{w}$ represents the quantization parameters for weights (clipping bounds $\alpha$ and $\beta$), and $\mathcal{D}$ is the training data distribution with signal samples $\mathbf{s}$ and labels $\mathbf{y}$. Using the fake quantization operator $F_{b}(\cdot)$ defined in (3), the optimization problem becomes:
@@
Note that in this work, activations remain in full precision throughout the network, allowing the neural receiver to preserve signal fidelity while achieving computational efficiency through weight quantization.
@@ 4 Experimental Results and Discussion
LoS comparisons. The lower panel of Figure 1 illustrates the BLER performance under LoS conditions using the CDL-D channel model. The 8-bit QAT configuration achieves 10% BLER within an SNR range of 3.94–3.97 dB across mobility scenarios, respectively, outperforming the LS baseline by approximately 2.4–4.8 dB as mobility increases. At velocities ranging from 10 to 35 m/s, QAT 4-bit requires 4.20–4.23 dB to attain 10% BLER, maintaining a performance advantage over LS of approximately 2.19–4.50 dB, depending on the mobility level. Results show that quantization-aware training is essential for reliable performance at ultra-low bitwidths, as the 4-bit QAT neural receiver consistently outperforms the PTQ variant across all UE speeds.
@@
5 Conclusion and Future Work
The 4-bit QAT meets 10% BLER targets in challenging NLoS scenarios where 4-bit PTQ fails, maintaining $\leq$0.8 dB SNR gap to FP32. We intentionally kept channel and network configurations identical to PTQ to enforce a controlled comparison and attribute gains specifically to the training methodology. Recent work shows that pruning and quantization trade off accuracy differently [7], having established a QAT baseline here, incorporating further compression methods via joint sparsity remains future work. Furthermore, our QAT framework can be extended to low-bit activation quantization and mixed-precision weight–activation designs, which we also leave for future work.
The $8\times$ weight compression enables hardware-software codesign: on-chip SRAM and bandwidth can target 4-bit storage while maintaining near-FP32 performance. This enables lightweight inference on resource-constrained edge devices, reducing memory and latency for 6G deployment. The strong LoS performance of 4-bit QAT receivers also positions them as viable candidates for integrated sensing and communication as well as site-specific deployments.